# A Research Roadmap for Augmenting Software Engineering Processes and Software Products with Generative AI

### v1

@@ 9. Threats to Validity
Construct validity concerns whether the key concepts investigated were correctly defined and operationalized. A potential threat lies in the subjective interpretation of what constitutes GenAI augmentation in software engineering processes and artifacts. To address this, the fundamental forms of GenAI in SE were initially elicited through collective discussions during the FSE 2025 “2030 Software Engineering” workshop and grounded in existing taxonomies, such as the one proposed by SEI. These constructs were subsequently refined and validated through multiple iterations of internal discussions and external feedback sessions involving peers. This triangulation of sources and perspectives aimed to ensure that the constructs reflect a shared and well-grounded understanding rather than the isolated view of a subset of authors. Another possible threat concerns the coverage of the Rapid Literature Reviews (RLRs) used in Design Cycle 2. The search strings adopted in each review may not have captured all relevant literature, potentially limiting the breadth of the investigated evidence. Moreover, the use of LLM-based filtering might have inadvertently excluded relevant publications. To mitigate this, samples of the excluded papers were manually inspected to validate the filtering process and verify that no relevant sources were systematically missed.
@@ 10. Conclusion and Perspectives
Based on our findings, we close with a forward looking perspective by offering ten predictions for software engineering in the year 2030:

### v2

@@ 9. Threats to Validity
Construct validity concerns whether the key concepts investigated were correctly defined and operationalized. A potential threat lies in the subjective interpretation of what constitutes GenAI augmentation in software engineering processes and artifacts. To address this, the fundamental forms of GenAI in SE were initially elicited through collective discussions during the FSE 2025 “2030 Software Engineering” workshop and grounded in existing taxonomies, such as the one proposed by SEI. These constructs were subsequently refined and validated through multiple iterations of internal discussions and external feedback sessions involving peers. This triangulation of sources and perspectives aimed to ensure that the constructs reflect a shared and well-grounded understanding rather than the isolated view of a subset of authors.
Another potential threat concerns the coverage of the Rapid Literature Reviews (RLRs) used in Design Cycle 2. The search strings adopted in each review may not have captured all relevant literature, potentially limiting the breadth of the investigated evidence. Moreover, the use of LLM-based filtering (using Qwen3 and Gemini-2.5 respectively) might have inadvertently excluded relevant publications due to the potential risks of bias and hallucination. To mitigate this validity risk, we manually inspected samples of the excluded papers to double check that no relevant literature was systematically missed. We also provide open access to the data of our literature reviews (person), 2026), thereby providing additional transparency and replicability of the RLR process.
@@ 10. Conclusion and Perspectives
We like to close by offering our ten predictions for software engineering in the year 2030: